Displaying reports 541-560 of 1091.Go to page Start 24 25 26 27 28 29 30 31 32 End
Reports until 16:03, Tuesday 23 July 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 16:03, Tuesday 23 July 2024 (11770)
Assorted notes from todays lab work

-Put the PDA36A2/PDA10A2 (TRANS/REFL for 775 respectively) on the table. Powered them with a new strip above this side of the table.

-Strung up 30+ ft homemade BNC cables for these signals (labelled Etienne/Achane).

-Moved the "floating" NIR camera to the other side of the room, to be used for 775 Transmission for the time being. If cameras are needed come talk to me. We should get more NIR cameras I think.

-Made several space saving optics mounts/base changes to more easily fit on the very cramped FC table.

-Prepping initial alignment into the cavity (still with no mirrors). Want to clean the PLA 3D printed alignment tool Daniel made. Seems like Iso is okay for this. Tested it on a reject print, small amount of Iso on a chem wipe, wipe clean, dry it. Waited a few minutes. No visible deformation on the test part so cleaning the real one now.

-Cleaned 3D printed alignment tool. Spent a fair amount of time using the two 775 steering mirrors to align into the first two OFC2 mirror positions (again no super optics on the table yet). With the super optics (tomorrow?) in place I believe the 775 light should fall into alignment fairly fast, these align tools have worked well in the past.

-Same thing as above but the 1550 light. Input should be first pass aligned.

-Aligned the PD and floating camera to the 775 first pass Trans path.

-Roughly aligned the flip up mirror in the 1550 Trans path onto the 50:50 BS for used later. 

-End of work particle count: 0/0/0

This is about everything I can think of to prep the second filter cavity. Will hopefully unbox 4 new super optics tomorrow.

Electronics General
Lee McCuller - posted 08:52, Tuesday 23 July 2024 (11768)
1U and 2U chassis to use for electronics assemblies in B111 makerspace

I moved the chassis I bought into the makerspace for use. NOTE: they should have a bag of anodized #4 screws somewhere with them. They have holes for these screws, but I found that they need to be tapped for the screws to work well. The anodization is too hard and thick for the screws to self-tap.

see notes and purchase

Please machine these as needed to mount electronics within.

I haven't moved them yet, but we also have D9 and D25 dsub hole punches. We are getting a BNC 0.5" D-shape punch, but it has a long lead. For now, just drill 0.5" holes and tighten washer if you need to add BNCs to front panels.

Atoms General
Briana Chen - posted 08:34, Tuesday 23 July 2024 (11764)
Collimation

[Briana, Ian]

Adjusted the fiber collimator by turning (first had to unlock it with a small screwdriver) to produce as collimated as possible of a beam. This is the current profile after adjustment: current780laserprofile- it is reasonably "collimated" for the scale of our distances, and we can put lenses to reshape it how we need. When using the beam profiler for this, the output power of the laser was 2.03 mW (current at 7.781 mA, temperature at 10.128 kOhms). 

Aligned for the new setup for measuring (tempsetup.svg) which gives room for beam profiling and placing lenses. Also, it will be easy to put in the EOM into this setup for the locking scheme without the pump beam. Cleared off some optics- unused optics are placed towards the middle of the table (IMG_2348.jpgIMG_2347.jpg). 

Measured immediately after laser and immediately after to verify that turning the waveplate changes the power reading. The values were respectively (for different laser powers), 7.30 mW -> 6.8 mW, 9.7 mW -> 8.93 mW, and 12.5 mW -> 10.9 mW, corresponding to ~93%,  92%, and 87% transmittance. Should verify this transmittance by also measuring the power after the waveplate and the reflectance to ensure losses are reasonable and within the beamsplitter specs. 

Note: for the damage threshold of the blue Nanoscan beam profiler, refer to this image from the specs: beamprofiledamagethreshold.png.

Images attached to this report
GQuEST SHG
Daniel Grass - posted 18:38, Monday 22 July 2024 (11765)
Fixed AOM Mount

[Erin, Daniel]

Since the Quanta Tech MT200 AOM to the Thorlabs PY005 adapter plates oriented the AOM at an undesirable orientation, I machined four #4-40 holes into the Isomet AOM to Thorlabs PY005 adapter plates. Erin cleaned the part, so it is ready for use in the lab. See attached files.

Non-image files attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 17:14, Monday 22 July 2024 (11762)
Profile and mode matching for output filter cavity 2

[Umran,Torrey]

We are profiling the two beams heading to OFC 2. Took a 60 second cumulative particle count that reads 1/0/0 at the start of the work. Random aside, we noticed on the new set of forks fully tightening the screw does not fully secure an optic. Something to look out for.

We first profile the beam coming out of the collimators, record these waists and waist locations, and use these as inputs for a JAMMT mode matching solutions. Current beam profiles for OFC2 can be found at "C:\Users\torre\Nextcloud\GQuEST\B102\Output Filter Cavities\OFC2\ModeMatching\* Beam profile.ipynb".

Then use JAMMT to calculate the mode matching of the second filter cavity. For now please note that I am using two 1/8 inch spacers to make OFC2 a total of 1/4 inch longer than OFC1. However, including and not including this quarter inch in my calculations for MM yields results that differ by amounts smaller than the precision in this process. JAMMT solutions can be found in the same directory, with screenshots.

I then implemented these solutions, note they differ by location by +/- inch, similar to other solutions. With the lenses in place I re-profiled (numbers in the same notebooks linked above). I can then plug these numbers into my finesse models for the second filter cavity. Attached are the results:775_OFC2_model.png 1550_OFC2_model.png. These calculation both expect having a mode mismatch of <1%. 

Besides alignment, the light should be ready for OFC2. Superoptics most likely tomorrow. End of work particle cound was 1/0/0. 

Other notes: The 1550 light is fairly astigmatic. We fiddled trying to fix this but am unsure where this is coming from. Long term this is a not a problem as this light wont be in the cavity anyways.

Images attached to this report
GQuEST End Mirror Mount
Daniel Grass - posted 14:35, Monday 22 July 2024 (11763)
Parts cleaned for End Mirror Mount

[Erin, Daniel]

Erin and I cleaned (baths, sonication, and scrubbing of 1:30 simple green:DI water, DI water, and isopropanol) and baked the custom parts from 3D Hubs and the helicoils at 150°C for 48 hours. We then cleaned the helicoil insertion and removal tools. We are ready to install the helicoils into the custom parts.

Atoms General
Briana Chen - posted 13:01, Monday 22 July 2024 - last comment - 20:01, Monday 26 August 2024(11760)
Absorption dip depth: effect from beam size through vapor cell

[Briana, Ian]

7/18: Started taking data with a 150 mm lens in but messed up wavelength in JamMt and alignment so data is void. However, from the initial look at that data, it seemed like putting in the lens removed the finer dips within larger dips (some of the absorption features are getting lost). However, wouldn't trust this too much because the alignment was bad. When I removed the lens, there was a dip in transmittance (blue circle) but by shifting the photodetector around, you can get back to a level close to the original with-lens measurement, although there is some noise (red circle), (see rip_alignment). This leads me to also think the issue in the last post could also be an alignment problem. However, I am not sure as I don't think I touched anything that would move it out of alignment, but who knows.

7/19: Realigned and measured for the 150 mm lens. First, I scanned across temperature with the lens in place. Then I removed the lens and scanned. Then I put the lens back in its original place, beam profiled the vapor cell, and moved the photodetector to see if I could get the same power level. I measured the beam profile on the active area of the photodetector, right before entering the vapor cell, and right after the vapor cell. The issue was the order in which I did this because it required a lot of putting the lens back to where it was, which might not have been accurate. A more consistent procedure is discussed below. At this point I also put in a 0.2 OD ND filter to not saturate the blue nanoscan beam profiler (didn't find a setting to increase saturation limit in the program).

7/21-7/22: Took data with the 100 mm lens (at 25 Celsius on 7/21, at 40 Celsius on 7/22). This is the procedure going forwards (can be repeated for the different focal lengths):

First, profile the beam. The laser power should not affect the beam profile but for consistency, set the laser current at 110 mA, temperature at 6.755 kOhms. Set the position of the lens and vapor cell (everything is relative to the mirror before the beam enters the vapor cell). Record the positions of where the vapor cell will begin/end, where the lens will be. Beam profile at the beginning of the vapor cell without the lens in place and record the location. Put the vapor cell in place. Beam profile after the vapor cell without the lens and record the location. Determine the location of the photodetector where the waist should be. Ensure that the beam profile you measure at the photodetector is within the 3.6 mm by 3.6 mm aperture and record the location/beam profile. Put photodetector down and mark the location. Slightly tune the positioning of the photodetector until you can see good dips (this was largely unsuccessful, I'm not sure how the tuning works). Do the scans/measurements (NO LENS). Put the lens in place. Confirm that the light falling onto the photodetector is the same as without the lens by checking alignment. Do scans (WITH LENS). Put in the beam profiler where the photodetector is (keeping the lens in place), which will require removing the photodetector. Then, beam profile after the vapor cell. Remove vapor cell and beam profile right before the lens.  

Using Torrey's beam profiling code and Jammt, you should be able to determine the beam size reasonably accurately. I was running into issues where the beam size I measured differed by >3 mm from JamMt because I should have picked 'thick lens' on the JamMt settings (not thin lens). With this fixed, there is a ~1mm difference in beam size than expected. Could be explained by error in distance measurements. Moving forward, I'll verify the distance measurements more accurately to see if this is the issue. Because I was having trouble with JamMt, I had started profiling the beam in the sequential manner, so I think the values for the beam size are correct, but just are not verifiable with JamMT.

Results:

The dip depth comparisons from all the data are shown here with the measured beam sizes (DipDepthComparison.png). Even at a different vapor cell temperature, the smaller beam size decreases the strength of the absorption dip (the 150 mm lens produces a smaller beam size by about ~1 mm). I want to justify this by saying that a smaller beam size means photons will interact with a lower number of atoms so less photon absorption occurs (in contrast, a larger beam means photons are more likely to be absorbed because they interact with more atoms). I feel like you would be saturating the atoms by bombarding them with photons, but I'm not sure if there are too many atoms for this to occur (spontaneous emission process is too fast?). Something I'll look into/calculate.

Also, if you plot the relative dip depth (1 - background voltage / signal voltage), you see that it decreases as you increase power, which would make sense as the increase in absolute dip depth is not enough to overcome the larger increase in background power (relative_dip_depth). This could also explain the saturation effect that we see (dip depth becomes less pronounced once you increase power too much). 

From the 7/19 plots, it seems like there is a power limit before the dip depth begins to decrease, which is due to some saturation effect. I want to say that this is a saturation limit on the photodetector, but as Ian pointed out it could be an issue with the Moku, so I should measure the power before reaching the photodetector to confirm. Based on the manual, if the photodetector is reading over 5V for a 50 Ohm load (which is what the Moku is set to), then it will saturate, but the output has consistently exceeded 5V output reading and has still retained an increasing dip depth. Should double check settings.

Smaller absorption features (on both dips) are still getting lost. If you get the right alignment, you can see the dips. Maybe the photodetector needs to be a certain distance away to register it but I'm not sure why/how since all the light is getting concentrated onto the active area anyways.

The way I measure the dip depth is to take the ends of the temperature vs. reading plot and fit an exponential to it (because the shape of the temperature curve sometimes trails up towards higher temperatures and is not a linear background). After subtracting it, I get something like the second slide. subtracting_background_example.pdf. Before I was doing it by hardcoding indices, so this should be better and easier to verify. 

See picture of the setup for the measurements. I plan to do the same with a 75 mm lens for a more drastic difference in beam size. Before that, the laser output will be collimated using the fiber collimators at the point of the alser. Ideally, light exiting the laser would be collimated. This makes it easier to mode match the beam, which we will need to do eventually for the probe/pump. 

Miscellaneous:

Connected CTL OUT of TED200C (laser temperature control) to input 3 of Moku (labelled Paige Bueckers). Must set the CTL OUT impedance of at least 10 kOhms from the TED200C manual, otherwise it will not register so on the Moku an impedance of 1 MOhm was set for the channel of CTL OUT input. This output is a voltage proportional to the actual temperature. The voltage range is 0 to 10 V with the conversion coefficient of 2 kOhms/V (using TH 20 K sensor). Using this signal instead of the previous triangular wave signal, which is more reflective of the actual temperature, did not solve the shifting issue: the frequency still has to be low to produce a not obvious shift. I thought this could be because of cable length differences between the CTL OUT BNC cable and the photodetector BNC cable, but probably not since electricity travels extremely fast. I also think that the temperature controller has asymmetry since cooling the laser may not have the same efficiency as heating it (this could be something with the PID controller of the laser temp controller).

Images attached to this report
Non-image files attached to this report
Comments related to this report
Briana Chen - 20:01, Monday 26 August 2024 (11767)

I saw this page and thought it was interesting how a hot spot on the laser looks kind of like an inverted absorption dip. I didn't know lasers could do this- it probably shows up in the beam profile. 

 

Environmental General
Lee McCuller - posted 11:33, Saturday 20 July 2024 (11707)
Lab layout (floorplan) and room numbers for new spaces.

Here is the layout an naming scheme for East Bridge new labs. It is up at https://nxc.mccullerlab.com/s/aEbdt4JeNPeen6k to be found in https://nxc.mccullerlab.com/apps/files/files/91750?dir=/Lab/Layouts

Non-image files attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 16:59, Friday 19 July 2024 (11761)
FC2/3 sled update

FC2/3 Sled is almost prepped for super optics. 1550 light is ready to be mode matched. The twice shifted via AOM 775 light is successfully fed to the FC2/3 sled via fiber. There is a 5 meter fiber for both wavelengths. On the sled a 2 m fiber for the 775 path and a 50:50 fiber BS for the 1550 path, as seen in this. The same will be done for 775 light for the third filter cavity. Tasks to be done:

-Mode match both wavelengths.

-Install PDs, wire up, etc.

-Assemble cavity optomechanics (flexture mounts)

-Superoptics.

Images attached to this report
Atoms General
Briana Chen - posted 23:40, Thursday 18 July 2024 (11745)
Absorption dip depth: effect from laser power/vapor cell temperature

[Briana, Ian]

We want to see how the laser power and the vapor cell temperature affect the strength of the absorption dip (the length of the dip). We want to first get as many atoms into the "ground state" for the 780 nm transition (5s 1/2) by heating the vapor cell. Then, we want to tune the laser power by changing laser current to excite as many atoms as possible into the 5P 3/2 transition state. This can maximize the strength of our absorption dip.

On 7/16: For each vapor cell temperature (25, 30, 40 Celsius), a temperature scan was taken for currents of 100, 110, 120, 125, 130, 135, and 140 mA. The sawtooth signal (100% symmetry) was used to scan across the laser temperature (which changes the wavelength output by the laser), but this caused some overlapping absorption dips due to the sudden signal drop causing the temperature to pass through the absorption dip again (this is apparent as seen here sawtooth_issues.pdf). So, this measurement was redone on 7/17 with a triangular wave, which has smooth continuous temperature shifts. We also measured the power output by the laser at each current/vapor cell temperature, data is here for both days of data. 

On 7/17: For each vapor cell temperature (25, 40 Celsius), a temperature scan was taken for currents of 110, 120, 125, 130, and 135 mA. A slower scan frequency (10 mHz) was used to ensure the temperature shift (shown to occur in the previous post) was not the same. To elaborate, from the last post there was a shift in the absorption dip along the temperature axis depending on whether the triangular wave signal at the absorption dip had a positive or negative derivative. This is most likely beause of response time delays, which cause slight asymmetry along the signal. At low frequencies the dips overlap properly but at higher frequencies (>50 mHz), shifting occurs (see whyisthereashift for results). Anyways, for some reason, the transmission started saturating and becoming noisy at 135 mA. The photodetector used to measure power was the same and measured 10.91 (on 7/16), 10.60 mW (on 7/17) for 135 mA at 25 degrees Celsius. I checked the specs of the photodetector and it can only read up to 5V for 50 Ohm impedance, which is what we are applying. We had no gain and the settings from the Moku were all the same, but I could be missing something. I also wonder if the photodetector was misaligned. 

Plotting the absolute dip strengths of these dips from both days shows an increasing absorption dip strength with vapor cell temperature (comparisons between vapor cell temperature (25, 30 and 40 Celsius) from 7/16 and 7/17 data are here: dipdepth.pdf), suggesting we are starting to pack the atoms into the S 1/2 state. From those graphs, the maximum absolute dip depth occurs around 125-130 mA (corresponding to 9-10 mW of laser power, as measured by the red photodetector). I think it makes sense to plot the absolute dip depth since that informs you of the amount of photons absorbed. As pointed out during the meeting, it's a bit strange that the absolute dip depth decreases upon saturation (would expect it to plateau), but I think this can be attributed to the detector being unable to distinguish changes in intensity at some limit, causing noise and loss of resolution.

Plots from 7/16 data on the absorption dips with temperature on the x axis are here 7_16_24temperature.pdf). Based on our laser spec sheet data (full excel file here, scanned datasheet here), increasing laser current or laser temperature increases laser wavelength. So, the absorption dip occurs at lower temperatures as you increase the current, which can be explained by the wavelength increase from current increasing being offset by the decreased temperature (corresponding to decreasing wavelength).

When we plot these temperature axis graphs in wavelength space (wavelength_absorption.pdf), we confirm that the values are reasonably close to the 780 nm transition. Offsets and patterns in the location of absorption are because of systematic calibration errors since conversion from temperature to wavelength was done through linear interpolation with the laser specs. This is all under the assumption that the output of the laser matches specs accurately. From our own measurements, there is obviously offset in power/temperature so overall these numbers are a rough estimate. 

Daniel measured a splitting of 6.8 GHz using graphs in the last post and I measured a splitting of 6.3 GHz from one dataset (from 7/17, with vapor cell of 25 deg Celsius). From the very initial estimate, I would guess that the transitions correspond to the F=1 (S_1/2 orbital) to the F=0 (P_3/2 orbital) and F=2 (S_1/2 orbital) to the F=3 (P_3/2 orbital) for the main and smaller dip, respectively. I calculated this by finding the average of the wavelength difference between the small dip and big dip. Then, I subtracted different possible hyperfine transitions between the S 1/2 and P 3/2 states (with selection rules being that deltaF = 0 or +/- 1) to see which one best matches the relative wavelength difference from data. This one dataset just happened to correspond to the above guess. A notebook showing these calculations is attached here: Hyperfineenergies.ipynb. Rb transitions found used found here (specific image used is rb_transition.png). Code will be uploaded to Gitlab. All data is in NextCloud under Users/briana.

Used oscilliscope mode to take data because data logger took too long for too little reward (files too large and low resolution). Assembled the other photodetector. We've been using the one labelled McCuller Lab for these measurements. We also assembled the now free-space EOM (see attached datasheet). We will need other lenses to focus the beam onto the 2mm aperture of the EOM (see updated schematic). 

Images attached to this report
Non-image files attached to this report
General
Torrey Cullen - posted 16:48, Thursday 18 July 2024 (11759)
GTrace Propagation of a beam for GQuest

Was looking over this post from a CE talk. This was mainly for fun to test using this code because I hadn't tried it yet. I made a filter cavity using the mirror specs and cavity dimensions. fullview.png is the result. zoomed_in_on_mirror.png is also kind of interesting on how the coating and mirror are treated. Not sure how useful this kind of image would be of the entire GQuEST experiment but I am confident I could built it if we wanted it.

Images attached to this report
Non-image files attached to this report
GQuEST SHG
Torrey Cullen - posted 16:28, Thursday 18 July 2024 (11758)
1st order twice shifted AOM light

Continuation of my previous post.The second AOM on the SHG sled now has light aligned so that the second pass, 1st order shift is easily accessable. Will align into a fiber and give the second filter cavity light tomorrow. Power budget is as follows:

-1st pass input to the AOM: 2.05 mW

-1st pass 1st order diffraction power: 1.7 mW

-2nd pass input to AOM: 1.7 mW

-2nd pass total power that reflects on the PBS: 1.3 mW (innefficiency from the quarter waveplate and PBS)

-2nd pass 1st order diffraction power: .6 mW

The second pass is much less efficient but this is still more power than we need to lock the cavity.

Another problem I encounter is the machine shopped metal adapter that attached the AOM to the PY005 5 axis stage is configured so that the longest side of the PY005 is along the beam propagation axis. This makes the space very limited on the SHG sled. I've found a work around, seen in the photos. Not the prettiest but it works. Long term we may want to redo these.

 

Images attached to this report
Lab Infrastructure General
Daniel Grass - posted 12:04, Thursday 18 July 2024 (11757)
Metal Containers and some Glassware Cleaned in Chemistry Room

I put the corresponding solvent into the metal bins labeled (SG + DI) (simple green (30:1) and DI water), DI (DI water), and Iso (isopropanol) and placed all 3 bins in the sonicator at once. The bins fit nicely into it. There was no visible degradation to the bins after this procedure. I then emptied them.

I noticed the sonicator temperature rose as soon as I turned it on. Considering this rapid change, I think this is an error with the sensor.

I also cleaned the 50 mL and 1 L beaker with simple green and then DI water.

I think the chemistry room is ready to use!

Atoms General
Daniel Grass - posted 01:31, Thursday 18 July 2024 (11756)
RbQ Final Vacuum Design and 3D MOT Photos

See attached JPGs. JPGs were much smaller than the PNG files for similar quantity. Let me know if you need PNGs.

Images attached to this report
Lab Infrastructure General
Daniel Grass - posted 21:33, Wednesday 17 July 2024 (11755)
Chemistry Room Cleaned and 1 Step Away from Being Ready to Clean Parts

I cleaned the chemistry room with some isopropanol and kim wipes and the shop vac (using the shop vac after chemicals are introduced is a worse idea so I wanted to start clean)

I made a standard operating procedure for cleaning parts. The general idea is these are parts that have no assumed cleanliness.

I think we should clean the containers that will hold the parts and then we will be ready.

GQuEST SHG
Torrey Cullen - posted 16:11, Wednesday 17 July 2024 (11754)
AOM 2 on SHG sled

[Erin, Torrey]

The homeade power cable for the RF power amplifier properly draws current. We are supplying the amplifier with the same variable power source used on the other AOM set up. We aligned the AOM with the power off. The power to the RF amplifier is from Moku 1. The output is at its max 2 Vpp with a 10 db attenuator on the end, limiting it to .63 Vpp input on the RF Amplifier. The amplifier applies a 29 +/- 1 dbm boost, so at most we have about 1 W of power going to the AOM. Note that the manual says 2.2 W will start to cause damage.The output of the moku can go above 250 MHz which is the max we should need for these AOMs. Note that the RF signal from Moku 1 channel 4 should never be put in +14 db mode. This WILL cause damage to the AOM. It might be smarter to operate it in +14 db mode and use an appropriate attenuator so that damage can never be caused.

With the AOM properly powered we observe a 1st and 2nd order beam. With 2.05 mw incident power we are getting 1.45 mw in the first order beam (71% efficiency) and a negligible amount in the 2nd order beam (10 uW ish). This will be more than we need to lock the second filter cavity, and may even need to be attenuated. Tomorrow we will attempt to reflect it back and get a second pass, first order beam in the fiber coupler. For now we will use a flat mirror, the curved mirror will be used on a future set up.

Lab Infrastructure General (GQuEST, Layouts)
Sander Vermeulen - posted 12:05, Wednesday 17 July 2024 (11752)
East Bridge Lab layout

Attached is the current version of the layout mockup for the new lab spaces.

Images attached to this report
Non-image files attached to this report
Lab Infrastructure General
Daniel Grass - posted 19:18, Tuesday 16 July 2024 (11751)
Chemicals moved to Chem Room (B110)

I moved the chemical we will use from B102A to B110 and placed them under or in the fume hood. Please see this list and update it accordingly if you move something.

There are still some chemicals left in B102A. There are four 1L bottles of Methanol and some solids like Cesium.

GQuEST SHG
Torrey Cullen - posted 16:27, Tuesday 16 July 2024 (11749)
AOM 1 out of cabinet on SHG sled

[Erin, Torrey]

I've pulled one AOM out of the 5 contained in this post. This is serial number 1002 and is labelled as such, as seen in the photo.

The max power on each of the four AOM paths on the SHG sled currently is <5 mW of 775 nm light. The spec sheet says the max power density of these AOMs is 5 W/mm^2. Our current solution has a waist of 57 um. This gives an optimistic max power density of 5e-3 W / (57 e-6)^ / (1m/1000mm)^2 = .64 W/mm^2. At the current power levels and this design, we should not be able to damage the AOM (optically at least). The beam diameter is supposed to be in between 200 and 300 um, therefore I will keep it as close to the lens as possible.

I am waiting to meet with Chris Wipf to ask if I can unplug our 4th moku from the cryo experiment, it seems fairly integrated. Going to use our fourth moku to power the AOM RF and lock OFC2. Erin made a power connection onto the ZHL-1-2W+ RF power amplifier since it does not come with one. We will similarly operate at max voltage on the moku with an attenuator at the input of the amplifier to prevent larger than intended signal amplitudes.

Images attached to this report
Atoms General
Briana Chen - posted 08:01, Tuesday 16 July 2024 - last comment - 19:14, Tuesday 16 July 2024(11743)
More absorption dips/temperature control setup

[Briana, Ian]

Set up TC300 (vapor cell temperature controller), which is next to the laser controllers. Even though the manual for the TC300 vapor cell temperature controller states that using a 4-wire setting for better accuracy, it does not apply to our thermistor. We should be on the 2-wire setting. If the control isn't applying, make sure the channel buttons are illuminated in green as otherwise they're disabled. Currently having an issue with the discretization of analog temperature output from TC300 as there is only a sharp dip in voltage (corresponding to temperature) when the temperature read reaches an integer (for example, from 23.0 to 22.01 Celsius, the voltage reading is roughly constant and once 22.0 Celsius is reached, the voltage reading drops significantly). See attached picture (temp_ctrl_discrete). I wonder if the 20 Ohm impedance from the analog output needs to be matched by the impedance BNC cable connecting to it. I will also try expanding the temperature range as the voltage drop is equal to the 5V spanned by the analog output divided by the set temperature range (0-50 Celsius), so maybe that has an effect. This range was chosen because the heater assembly should not exceed 50 deg Celsius. Either way, will need to fix this before tuning PID controllers of the TC300. 

We moved the bottom Moku Pro (not connected to anything) on the rack near the main computer in place of where the Moku Go was. Moku Go has been returned to original cabinet. Both channels of the vapor cell temperature controller (labelled Aliyah Boston and Angel Reese) have been connected to the Moku input. Photodetector connected to Moku input is labelled Caitlin Clark. Moku output (connected to laser temperature controller input) is labelled Joey Votto. 

We repeated the measurements last post with the current controller set. We set the laser current controller at 125.08 mA, laser temperature controller scanning from 6.871 - 7.160 kOhms (32.8 - 33.78 Celsius), and the temperature setting on the vapor cell to 45 degrees Celsius on both controllers. Current limit is 0.7 A. We see two dips, one small and one large. The large one likely corresponds to the 780 nm transition, but maybe there is some hyperfine structure causing the smaller dip. We can determine the temperature at which these peaks occur and then, using laser specs, the wavelength, which we can then use to deduce the transition (need to call Thorlabs to ask for the data since it's specific to our laser). I have not figured out why there is a shift between the trials- it corresponds to the dips on the signal downfall (groups 2 and 4) but not sure why since the signal is the same at these points. If you zoom in onto the smaller dip, you see two overlapping profiles, potentially from two closely spaced transitions. Will determine what wavelengths these correspond to after getting Thorlabs data. Also, previously when we measured the absorption, we had a laser current of around 203 mA, so it was probably saturating the photodetector.

We mounted and placed the EOM (EO-PM-NR-C1) on the vapor cell setup board. Attached is the scanned sheet it came with and the new schematic since the EOM is now free space (need an additional 780 mirror). We will need lenses for focusing the beam down to the 2mm aperture and possibly a waveplate to make sure the input beam is vertically polarized. Also learned to make BNC cables (https://wiki.mccullerlab.com/DCC/T2400001). 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Daniel Grass - 19:14, Tuesday 16 July 2024 (11750)

Based on the laser specs found here, the spacing between the peaks matches the Rb87 data found here to ~10%.

Displaying reports 541-560 of 1091.Go to page Start 24 25 26 27 28 29 30 31 32 End